# Wave-packet dynamics in a graphene with periodic potentials

### v1

@@ IV Conclusions
In this work, we used the Dirac continuum model to study the propagation of electronic wave packets in graphene with periodically arranged potential barriers. Our simulations reveal that both the size and arrangement of the scattering centers significantly affect transmission probabilities. Increasing the barrier height leads to non-monotonic changes in transmission, with larger wave packets being more impacted. The arrangement of potential barriers, whether symmetric or staggered, plays a crucial role, with staggered configurations reducing transmission more effectively. These findings enhance our understanding of wave packet dynamics in graphene and could inform the design of graphene-based electronic devices.

### v2

@@ IV Conclusions
In this work, we used the Dirac continuum model to study the propagation of wave packets representing low-energy electrons in monolayer graphene with periodically arranged potential barriers. The system studied here provides insights into the effects of spatially varying electrostatic potentials originating from positive and negative charge puddles commonly found in real graphene samples. Our simulations reveal that both the size and arrangement of scattering centers significantly influence transmission probabilities. Increasing the barrier height leads to non-monotonic changes in transmission, with larger wave packets being more strongly affected. The configuration of potential barriers – whether symmetric or staggered – also plays a crucial role, with staggered arrangements leading to greater transmission suppression. These findings deepen our understanding of wave packet dynamics in graphene and offer useful guidance for the design of future graphene-based electronic devices.
@@ V Acknowledgements
AC acknowledges financial support from the Brazilian National Council for Scientific and Technological Development (CNPq), through the UNIVERSAL and PQ programs, and from the Brazilian Coordination for the Improvement of Higher Education Personnel (CAPES), through the PROBRAL program.